nettym:

Match the following options to question. (ASTRONOMY) a. Force would go up by a factor of 2. b. Force would go up by a factor of 4. c. Force would go down by a factor of 2. d. Force would go down by a factor of 4. Consider the gravitational force between the Earth and the Moon. If you were to double the mass of the Moon, how would the gravitational force between them change? If you were to double the mass of both the Earth and Moon, how would the gravitational force between them change? If you were to decrease the distance between the Earth and Moon by a factor of 2, how would the gravitational force between them change? If you were to double both masses and increase the distance between the Earth and Moon by a factor of 2, how would the gravitational force between them change?

MAGABACK:

@toga @hero tbh maybe they know. I'm not good with physics

adrianluvvsyouu2:

The equation Force = G * m1 + m2/r^2 is helpful when considering gravitation problems in physics. When doubling mass of the moon it becomes F = G * m1 * (2*m2)/r^2, doubling m2 has the effect of doubling the force entirely, as all variables are being multiplied. If it makes more sense to write it as F = 2*G*m1*m2/r^2 that works as well Same idea for the second question, you replace m1 with 2*m1 and m2 with 2*m2, effectively quadrupling the force Decreasing the distance (r) by a factor of 2 just means division by 2. So in the formula, plug in (r/2) in the denominator. That entire quantity, when squared, gives F = Gm1m2/(r^2/4) When a fraction is on the bottom part of another fraction, its denominator is multiplied by the numerator Meaning a/(b/c) = c * a/b. The same concept applies here, meaning the 4 in the denominator of the force equation is multiplied by the entire fraction. Therefore, it increases by a factor of 4. The final question I leave to you to answer using this, have fun with physics!

ykbatman:

you could probably just look dat up